# On an approach to canonicalizing elliptic Feynman integrals

### v2

@@ 3.2.1 Definition
We need to introduce a complete elliptic integral of the third kind depending on $e_{6}$ for both sectors following Sec. 2.3. However, we have already known no such integral is needed in Jiang:2023jmk . We can find there are relations between complete elliptic integral of the first kind and the third kind [^13] One may find similar relations applying the same procedure to equal-mass sunrise family. It is conjectured that such relations are associated with the existence of ${\rm d}\log$-forms, where no introduction of elliptic integrals is needed, and we can find such ${\rm d}\log$-form here with a constant linear combination of integrals of the third kind and the first kind.
$\vartheta_{0}(e_{6},\lambda)=-\frac{\varpi_{0}(\lambda)}{1+\sqrt{1-16y_{s}}}\,,\quad\vartheta_{1}(e_{6},\lambda)=-\frac{\varpi_{1}(\lambda)}{1+\sqrt{1-16y_{s}}}-1\,,$ (86)
and their explicit forms can be simplified as [^14] We have used special relations in App. D, the relations can also be used to express the periods $\varpi_{0}(\lambda)$ and $\vartheta_{0}(e_{5},\lambda)$ with simpler arguments. However, since the explicit expressions for the periods are immaterial for our discussions, we do not present them here.
$\Pi(k,k^{2})=\frac{1}{2}K(k^{2})+\frac{\pi}{4(1-k)}\,,\quad\Pi(1-k,1-k^{2})=\frac{(1+k)K(1-k^{2})}{2k}\,.$ (87)
Similarly, branch points ${c_{1},c_{2},c_{3},c_{4}}$ for the electroweak double box family are given by [^15] Here we define $z=-z_{9}+z_{6}+c_{5}$. Besides, there are two distinct elliptic curves for the electroweak double box family, and here, we only consider elliptic sectors associated with the elliptic curve for the top sector. The canonical basis for the whole family is given in Schwanemann:2024kbg .
@@
where we define new variables $x$ and $y$ to simplify the expressions
@@
The variables relevant in the Legendre norm form are
@@
One can deduce there must exist special relations like Eqs. (86) and (87) for the same reason, which are
@@
We will mainly study the behaviors of the integrals in the vicinity of the degenerate point $m^{2}\to 0^{-}$ (as in Jiang:2023jmk ), where the order of the branch points is $c_{1}<c_{2}<c_{3}<c_{4}$ and we have $y_{s}\to 0$, $\lambda\to 1$. The vicinity of the degenerate point $s\to-16m^{2}$ (corresponding to $y_{s}\to 1/16$ and $\lambda\to 0$) can be studied similarly. For the study of the point $s\to 0$, it is better to modify the normalization of the integrals in Eq. (1) by setting $\mu^{2}=m^{2}$.
Note that the shape of the elliptic curve and the integrals in the triangle sub-sector only depend on the kinematic variable $y_{s}$. This is why the maximal cut of the triangle integrals can be characterized by modular forms. The integrals in the top sector, on the other hand, depend on the variable $y_{t}$ as well. This leads to extra complications in the construction of canonical bases. In the next two subsubsections, we first consider the construction of a pre-canonical basis under the elliptic (triangle) cut with nice properties, then apply the transformation ${\cal T}_{\rm can}^{(\varepsilon)}$ to remove residual non-$\varepsilon$-factorized elements, and finally transform the differential equations to a canonical form with ${\cal T}_{\rm can}^{(0)}$.
3.2.2 Transformation for the elliptic sectors
We note that the elliptic cut of $I_{111110100}$ can be written as
${\rm EllCut}[I_{111110100}]=\int_{\cal C}\,\bar{u}(z_{6},z)\frac{{\rm d}z_{6}}{\sqrt{P_{2}(z_{6},z)}}\wedge\frac{{\rm d}z}{\sqrt{P_{4}(z)}}=-\int_{\cal C}\,\bar{u}(z_{6},z){\rm d}\log\left(\frac{1+\sqrt{\frac{z_{6}-c_{6}}{z_{6}-c_{7}}}}{1-\sqrt{\frac{z_{6}-c_{6}}{z_{6}-c_{7}}}}\right)\wedge\frac{{\rm d}z}{\sqrt{P_{4}(z)}}\,,$ (95)
where $\bar{u}(z_{6},z)$ is the reduced twist
Similarly, we have
$=\int_{\cal C}\,\bar{u}(z_{6},z){\rm d}\log\left(\frac{1+\sqrt{\frac{c_{7}(z_{6}-c_{6})}{c_{6}(z_{6}-c_{7})}}}{1-\sqrt{\frac{c_{7}(z_{6}-c_{6})}{c_{6}(z_{6}-c_{7})}}}\right)\wedge\frac{{\rm d}z}{(z-c_{5})\sqrt{P_{4}(z)}}\,.$ (98)
From Eqs. (95) and (97), we can see explicitly that they share the same elliptic curve defined by $w^{2}=P_{4}(z)$, and a natural idea is to construct the pre-canonical basis integrals as ${\rm d}\log\wedge\phi$.
Following the construction in Sec. 2, the four pre-canonical basis integrals in the top sector can be defined with
$\phi_{1}^{\rm npdb}$ $=\frac{1}{2}\left(1+\sqrt{1-16y_{s}}\right)\frac{{\rm d}z}{\sqrt{z(z-1)[z(z-1)+4y_{s}]}}\,,$ (99a)
$\phi_{2}^{\rm npdb}$ $=\frac{z\sqrt{(y_{t}-1)y_{t}[4y_{s}+(y_{t}-1)y_{t}]}}{y_{t}(z-y_{t})}\frac{{\rm d}z}{\sqrt{z(z-1)[z(z-1)+4y_{s}]}}\,,$ (99b)
$\phi_{3}^{\rm npdb}$ $=-z\frac{{\rm d}z}{\sqrt{z(z-1)[z(z-1)+4y_{s}]}}\,,$ (99c)
$\phi_{4}^{\rm npdb}$ $=\frac{z-\frac{1}{2}\left(1-\sqrt{1-16y_{s}}\right)}{z}\frac{{\rm d}z}{\sqrt{z(z-1)[z(z-1)+4y_{s}]}}\,,$ (99d)
for electroweak double box
$\phi_{1}^{\rm ewdb}$ $=(1+x)\sqrt{(3-x)(1+x)}\frac{{\rm d}z}{\sqrt{z(z-1)[4z-(1-x)^{2}][4z-(1+x)^{2}]}}\,,$ (100a)
$\phi_{2}^{\rm ewdb}$ $=\frac{z\sqrt{y(y-1)[x^{4}-2(1+4y)x^{2}+(1-4y)^{2}]}}{y(z-y)}\frac{{\rm d}z}{\sqrt{z(z-1)[4z-(1-x)^{2}][4z-(1+x)^{2}]}}\,,$ (100b)
$\phi_{3}^{\rm ewdb}$ $=-4z\frac{{\rm d}z}{\sqrt{z(z-1)[4z-(1-x)^{2}][4z-(1+x)^{2}]}}\,,$ (100c)
$\phi_{4}^{\rm ewdb}$ $=\frac{\sqrt{(3-x)(1+x)}}{3-x}\frac{4z-x^{2}+2x-1}{z}\frac{{\rm d}z}{\sqrt{z(z-1)[4z-(1-x)^{2}][4z-(1+x)^{2}]}}\,.$ (100d)
For the triangle sector, we only need to pick the first and the last ones [^16] There are $7$ basis integrals in the elliptic family defined by Eq. (79), while we know there are only $6$ master integrals for the first family in fact. The reason is the symmetric relation, which cannot be seen by twisted cohomology. Conventionally, we can choose an element vanishing under the symmetric relation and discard information about it in the projection transformation to Feynman integrals.. Moreover, we find that if we further change the integral of the second kind, namely, $\phi^{\rm npdb}_{4}$, with the derivative of the first kind $\phi^{\rm npdb}_{1}$ with respect to $y_{s}$, and then apply the transformations to the canonical bases for top sector and triangle sector, the mixing is $\varepsilon$-factorized as well. We note that the derivative basis cannot promise the $\varepsilon$-factorization for the electroweak family, but it is very close to so that we can achieve it with a simple transformation.
We would like to give short comments on the procedure above. Here, we want to construct a canonical basis for the whole elliptic sectors instead of sector by sector or, say, block by block, respectively. Our strategy is to try constructing basis integrals in the form ${\rm d}\log\wedge\phi$. In this way, the corresponding projection transformation contains more information in the mixing between sectors, which cannot be seen from separate maximal cuts. Moreover, it turns out that no further transformation is necessary for integrals of the first kind and the third kind. Similar to the case under the maximal cut, the integral of the second kind is always the most complicated one and it seems that by assigning it with the derivative basis of the first kind integral, the mixing becomes simpler. Specially, for the first family, simply using the derivative with respect to $y_{s}$ is enough. A more natural choice is the derivative with respect to $\lambda$ (and the other variable is $e_{5}$), but it is not as simple as $y_{s}$ in this case. Further transformation is needed. The simplicity with respect to $y_{s}$ is, in fact, an accident since it is not enough for the electroweak cousin of this family. Another point is the additional transformation is only dependent on the first kind of the triangle family. However, the two cases are very similar, and thus, one needs to check further non-trivial examples for this observation.
3.2.3 Mixing with non-elliptic sectors
Here we go on to find the canonical basis for the first family without cuts. As mentioned before, it is sometimes simpler to work with a pre-canonical basis instead of the canonical basis since there are no elliptic integrals in a pre-canonical basis. If we manage to make the mixing with non-elliptic sub-sectors for integrals of the first and the third kind $\varepsilon$-factorized, the same structure will be preserved in canonical basis, with applying ${\cal T}_{\rm can}$ or ${\cal T}_{\rm can}^{(\varepsilon)}$. Thus, we only need to modify the integral of the second kind later. Usually, we need not introduce elliptic integrals to make a pre-canonical basis satisfy this property, which makes the work simpler. Here we would like to show an example.
Provided with the canonical basis integrals in the elliptic sectors, we restore top sector integrals back to pre-canonical ones and then the structure of the coefficient matrix $\bm{A}$ can be presented schematically as
[…]
where the cyan cells are already $\varepsilon$-factorized, the purple cells are linear in $\varepsilon$, and the red cells contain only $\varepsilon^{0}$ terms.
There is only one element ${\bm{A}}_{4,8}$ containing $\varepsilon^{0}$-term in the mixing for $I_{4}$ with non-elliptic sub-sectors, and we apply an ansatz transformation to remove it
$I_{1}\to I_{1}+f_{1,8}(y_{s},y_{t})M_{8}\,,\quad I_{4}\to I_{4}+f_{4,8}(y_{s},y_{t})M_{8}\,,$ (102)
where the solutions are given by
Now all the mixing elements ${\bm{A}}_{4,j}$’s for $(j=7,\cdots,36)$ are in the forms $a^{(1)}\varepsilon+a^{(2)}\varepsilon^{2}$, and ${\bm{A}}_{1,j}$’s are all $\varepsilon$-factorized while ${\bm{A}}_{2/3,j}$’s are in linear forms $a^{(0)}+a^{(1)}\varepsilon$. According to (101), all the $A_{m,2/3}$’s are $\varepsilon$-factorized, and thus we can apply ansatz transformations for $I_{2/3}$
$I_{2/3}\to I_{2/3}+f_{2/3,j}(y_{s},y_{t})M_{j}\,,$ (104)
with which we can remove ${\bm{A}}_{2/3,j}^{(0)}$, without altering the $\varepsilon$-factorized structures for $I_{1}$ and $I_{4}$.
Then, we only need to modify the integral of the second kind, namely, $I_{4}$. We apply transformation ${\cal T}_{\rm can}^{(\varepsilon)}$, where non-$\varepsilon$-factorized elements in the corresponding connection matrix for the top sector are strictly lower triangular. The mixing for the integral of the second kind are all in the linear forms $a^{(0)}+a^{(1)}\varepsilon$, and if we remove all the $\varepsilon^{0}$-terms, the connection matrix is promised to be canonical directly with a further ${\cal T}_{\rm can}^{(0)}$.
The $\varepsilon^{0}$-terms can be removed with regular “integrating out” strategy, and it turns out there are no new functions needed in the transformation beyond the periods $\varpi_{0}(\lambda)$ and $\vartheta_{0}(e_{5},\lambda)$ in ${\cal T}_{\rm can}$. For simplicity, we suppress the explicit expressions here, which can be found in ancillary files.
@@ 4 Conclusion and outlook
In this paper, we develop a systematic approach to constructing a canonical basis for elliptic Feynman integral families involving multiple kinematic scales, which exhibit a univariate integral in the Baikov representation under maximal cut. The main idea is to perform the construction using the Legendre normal form of elliptic curves, where the geometric information is explicit, and then apply a Möbius transformation to map it back to the original integral family. The pre-canonical basis integrals under maximal cut are expressed as linear combinations of Abelian differentials without elliptic integrals in the combination coefficients. This feature provides simplification when performing the $\varepsilon$ factorization for the dependence on sub-sectors. Furthermore, the canonical bases exhibit favorable behaviors around chosen degenerate points, where they asymptotically take the ${\rm d}\log$-forms. This helps to obtain simple boundary conditions for the differential equations. Around the degenerate point, only the dependence of the basis integrals on the marked points becomes significant, allowing the integrals to be expressed in terms of MPLs in this context.
@@
There can be different ways to construct the pre-canonical bases and the transformations bringing them to the canonical ones, and the approach we offer here is largely independent of the specific construction. Besides the method described in the main text, which is also equivalent to the one from Gorges:2023zgv , we have explored several other possibilities which are described in the appendices. In App. A, we show that our pre-canonical basis can be transformed to a canonical one by converting the period matrix into a constant matrix without $\varepsilon$-dependent transformation vonManteuffel:2017hms ; Frellesvig:2021hkr . Despite a simpler $\varepsilon$-dependence, this method introduces numerous elliptic integrals into the transformation, which, in turn, appear in the $\varepsilon$-factorized differential equations. In addition, this approach inherently leads to the transformation with logarithmic behaviors near degenerate points due to the complementary arguments $\lambda$ and $1-\lambda$ in elliptic integrals of the first and second kinds. Relevant discussions are included in Frellesvig:2023iwr . In contrast, the approach present in the main text avoids these issues, and the resulting canonical bases satisfy most of the nice properties advocated in Dlapa:2022wdu ; Frellesvig:2023iwr ; Duhr:2024uid .
@@
To demonstrate our method, we use fully massive sunrise families and non-planar double-box families with an internal mass as proof of concept. We emphasize again that, while our construction of the pre-canonical bases in the top-sectors and the transformation to canonical ones are equivalent to the methods in vonManteuffel:2017hms ; Adams:2018bsn ; Frellesvig:2021hkr ; Giroux:2022wav ; Dlapa:2022wdu ; Pogel:2022vat ; Gorges:2023zgv , these only need to be done once and for all, and the universal results have already been given in our paper. The applications to different integral families merely boil down to appropriate Möbius transformations. Moreover, since the proposed approach is not inherently limited to elliptic cases, it is expected to extend to more complex geometries, e.g., hyperelliptic cases, as explored in the recent work Duhr:2024uid . For example, using the canonical basis integrals constructed for the four-parameter Lauricella functions, one may derive the canonical basis for the equal-mass non-planar double (crossed) box family, at least under the maximal cut. One can also try constructing a canonical basis in the Rosenhain normal form family with kinematic marked points, which can be potentially applied for generic hyperelliptic Feynman integrals.
@@
Our method can be applied to integral families where multiple sectors involve the same elliptic curve. In such cases, the construction can be done sector-by-sector. However, we have found that it is helpful to perform the construction for these sectors together by loosening the maximal cut. This has been demonstrated for the non-planar double box family in Sec. 3.2, where both the top sector and the triangle sub-sector are elliptic. It remains possible that one may construct the full integrand without any cut, combining the elliptic-sector construction in this work and the ${\rm d}\log$-form construction in Chen:2020uyk ; Chen:2022lzr . It is also interesting to explore the application of our method to cases where multiple elliptic curves are involved. In this case, there can be non-trivial interplays among the Abelian differentials in different elliptic sectors, and it is interesting to see how these can be handled within our approach. Another complication arises in the boundary conditions since the degenerate points of the different elliptic curves may not coincide. It remains to be seen whether one may utilize the degenerate information separately, and connect them via the differential equations. We leave these for future investigations.

### v3

@@ 3.4.2 Mixing with non-elliptic sectors
where the cyan cells are already $\varepsilon$-factorized, the purple cells are linear in $\varepsilon$, and the red cells contain only $\varepsilon^{0}$-terms. We would like to add sub-sector components to $\vec{I}^{\text{pre}}$ such that the above structure is not spoiled, and we can then rotate it to the canonical basis using the transformation matrix $\mathcal{T}_{\text{can}}$ derived in Sec. 2.1.
We first deal with the $\varepsilon^{0}$-terms in the sub-sector dependence of $I_{4}^{\text{pre}}$, which only appear in the matrix element ${\bm{A}}_{4,8}$. We denote the canonical basis for the non-elliptic sub-sectors as $M_{j}$ for $j=7,\cdots,36$. Introducing two coefficient functions $f_{1,8}$ and $f_{4,8}$, we apply the following transformation to $I_{1}^{\text{pre}}$ and $I_{4}^{\text{pre}}$:
$I_{1}^{\text{pre}}\to I_{1}^{\text{pre}}+f_{1,8}(y_{s},y_{t})M_{8}\,,\quad I_{4}^{\text{pre}}\to I_{4}^{\text{pre}}+f_{4,8}(y_{s},y_{t})M_{8}\,.$ (98)
Requiring that the $\varepsilon^{0}$-terms vanish in the transformed differential equations, we can find the solution for the two functions:
After the above transformation, the matrix elements ${\bm{A}}_{1,j}$ for $j=7,\cdots,36$ are all $\varepsilon$-factorized, while ${\bm{A}}_{4,j}$ are in the form $a^{(1)}\varepsilon+a^{(2)}\varepsilon^{2}$. We now turn to the $\varepsilon^{0}$-terms in the sub-sector dependence of $I_{2}^{\text{pre}}$ and $I_{3}^{\text{pre}}$. According to (97), all the matrix elements $A_{i,2/3}$ are already $\varepsilon$-factorized. Therefore, we can apply the following transformations to $I_{2/3}^{\text{pre}}$:
$I_{2/3}^{\text{pre}}\to I_{2/3}^{\text{pre}}+f_{2/3,j}(y_{s},y_{t})M_{j}\,$, (100)
without altering the $\varepsilon$-factorized structure for $I_{1}^{\text{pre}}$ and $I_{4}^{\text{pre}}$. The functions $f_{2/3,j}(y_{s},y_{t})$ can be solved by requiring the vanishing of $\varepsilon^{0}$-terms. Their expressions are rather lengthy, and we choose to not present them here.
With the above preparation, we can now apply the transformation ${\cal T}_{\rm can}$ to convert $\vec{I}^{\text{pre}}$ into $\vec{I}^{\text{can}}$. From the structure of ${\cal T}_{\rm can}$, it is easy to see that the sub-sector dependence of $I^{\text{can}}_{1/2/3}$ is still $\varepsilon$-factorized. On the other hand, the $\varepsilon^{0}$-terms re-appear in the differential equations of $I^{\text{can}}_{4}$, which is now in the linear form $a^{(0)}+a^{(1)}\varepsilon$. These $\varepsilon^{0}$-terms can be removed by one-fold integrations as before. The final canonical basis for Feynman integrals without cuts is given in ancillary files.
3.5 Non-planar double box B
Figure 5: The diagram for the non-planar double box family B is shown above. It is similar to Fig. 4, with two internal lines becoming massless.
Our final example is the non-planar double box B family shown in Fig. 5. It shares a lot of similarities with the family in Sec. 3.4, and the canonical basis is studied in Schwanemann:2024kbg . The kinematics is the same as Eqs. (80) and Eqs. (81). The propagators are also the same as Eqs. (79) except that $\chi=0$, so that $D_{1}$ and $D_{3}$ are massless. For this family, we will only consider the top sector and the elliptic triangle sub-sector [^9] Note that there is actually another elliptic curve associated with a different sub-sector, which we do not consider here., which share the same elliptic curve. The starting master integrals are chosen as
$\big\{$ $I_{21111110(-1)},I_{11121110(-1)},I_{11111110(-1)},I_{111111100},I_{211110100},I_{111210100},I_{111110100}\big\}\,.$ (101)
The first $4$ integrals belong to the top sector, while the other $3$ belong to the triangle sub-sector. We have one more master integral in the triangle sector than the case in Sec. 3.4 because the diagram has fewer symmetry relations.
As the case in Sec. 3.4, we perform two-variable constructions in the form ${\rm d}\log\wedge\phi^{\rm can}$ under the elliptic triangle cut, where the mixing between the two sectors is taken into account. The elliptic cut of $I_{111111100}$, $I_{111110100}$ and the reduced twist shares the same form as Eqs. (90), (88) and (89), respectively. Defining $z=-z_{9}+z_{6}+c_{5}$, we can express the branch points $c_{1},c_{2},c_{3},c_{4}$} for this family as
@@ 3.5 Non-planar double box B
where we have introduced the variables $x$ and $y$ to simplify the expressions:
@@
After the Möbius transformation, the relevant variables are
@@
Inserting the variables into Eq. (10), we readily obtain the integrands for the canonical basis. As before, we may find special relations similar to Eqs. (95) and (96), which are
@@
Applying the derivative trick as in earlier examples, we find that there is still one entry $\tilde{{\bm{A}}}_{4,5}$ in the differential equations that are not $\varepsilon$-factorized. This can be removed by a simple transformation:
$I_{4}^{\rm can}\to I_{4}^{\rm can}-\frac{2x^{2}\varpi_{1}(\lambda)^{2}}{3(x-3)(x+1)^{3}}I_{5}^{\rm can}\,.$ (109)
Alternatively, we can also perform the transformation at the level of the pre-canonical basis:
$I_{4}^{\rm pre}\to I_{4}^{\rm pre}-\frac{16x^{2}\varepsilon}{3(x-3)(x+1)^{3}}I_{5}^{\rm pre}\,$, (110)
before acting the matrix $\mathcal{T}^{\rm can}$. Both methods lead to a canonical basis within the two sectors.
@@ 4 Conclusion and outlook
In this paper, we present an integrand-level canonical basis Eqs. (10) for elliptic Feynman integral families involving multiple kinematic scales, which exhibit a univariate integral in the Baikov representation under maximal cut. The applications to specific integral families merely boil down to plug the information of branch points into Eqs. (2.1) and Eqs. (10). Furthermore, the canonical bases exhibit favorable behaviors around chosen degenerate points, where they asymptotically take the ${\rm d}\log$-forms. This helps to obtain simple boundary conditions for the differential equations. Around the degenerate point, only the dependence of the basis integrals on the marked points becomes significant, allowing the integrals to be expressed in terms of MPLs in this context.
Our derivation of the canonical basis is to perform the construction using the Legendre normal form of elliptic curves, where the geometric information is explicit, and then apply a Möbius transformation to map it back to the original integral family, making the results applicable for generic families. The pre-canonical basis integrals under maximal cut are expressed as linear combinations of Abelian differentials without elliptic integrals in the combination coefficients. This feature provides simplification when performing the $\varepsilon$-factorization for the dependence on sub-sectors. We emphasize again that, while our construction of the pre-canonical bases in the top-sectors and the transformation to $\varepsilon$-factorized (canonical) ones are equivalent to the methods in vonManteuffel:2017hms ; Adams:2018bsn ; Frellesvig:2021hkr ; Giroux:2022wav ; Dlapa:2022wdu ; Pogel:2022vat ; Gorges:2023zgv , these only need to be done once and for all, and the universal results have already been summarized in Eqs. (10).
To demonstrate our method, we apply it to five integral families and obtain canonical basis with success, where two of them are new results in this paper. One of them is under the maximal cut in Sec. 3.2, and the other is for the full integral family without any cuts in Sec. 3.3. Our method can be applied to integral families where multiple sectors involve the same elliptic curve. In such cases, the construction can be done sector by sector. However, we have found that it is helpful to perform the construction for these sectors together by loosening the maximal cut. This has been demonstrated for the non-planar double box families in Sec. 3.4, Sec. 3.5, where both the top sector and the triangle sub-sector are elliptic.
@@
Our idea of using the Legendre normal form combined with a Möbius transformation to derive a canonical basis is quite generic. It is largely independent of the particular choice of the pre-canonical bases and the canonicalization transformations. The canonicalization transformation described in the main text is equivalent to the one from Gorges:2023zgv . We have explored several other possibilities which are described in the appendices. In App. A, we show that our pre-canonical basis can be transformed to an $\varepsilon$-factorized one by converting the period matrix into a constant matrix without $\varepsilon$-dependent transformation vonManteuffel:2017hms ; Frellesvig:2021hkr . Despite a simpler $\varepsilon$-dependence, this method introduces numerous elliptic integrals into the transformation, which, in turn, appear in the $\varepsilon$-factorized differential equations. In addition, this approach inherently leads to the transformation with logarithmic behaviors near degenerate points due to the complementary arguments $\lambda$ and $1-\lambda$ in elliptic integrals of the first and second kinds. Relevant discussions are included in Frellesvig:2023iwr . In contrast, the approach present in the main text avoids these issues, and the resulting canonical bases satisfy most of the nice properties advocated in Dlapa:2022wdu ; Frellesvig:2023iwr ; Duhr:2024uid .
@@
Moreover, since the proposed workflow is not inherently limited to elliptic cases, it is expected to extend to more complex geometries, e.g., hyperelliptic cases, as explored in the recent work Duhr:2024uid . For example, using the canonical basis integrals constructed for the four-parameter Lauricella functions, one may derive the canonical basis for the equal-mass non-planar double (crossed) box family, at least under the maximal cut. One can also try constructing a canonical basis in the Rosenhain normal form family with kinematic marked points, which can be potentially applied for generic hyperelliptic Feynman integrals.
@@
It remains possible to extend our two-variable construction in Sec. 3.4 and Sec. 3.5. One may construct the full integrand without any cut, combining the elliptic-sector construction in this work and the ${\rm d}\log$-form construction in Chen:2020uyk ; Chen:2022lzr . It is also interesting to explore the application of our method to cases where multiple elliptic curves are involved. In this case, there can be non-trivial interplays among the Abelian differentials in different elliptic sectors, and it is interesting to see how these can be handled within our approach. Another complication arises in the boundary conditions since the degenerate points of the different elliptic curves may not coincide. It remains to be seen whether one may utilize the degenerate information separately, and connect them via the differential equations. We leave these for future investigations.